# Towards On-Device Personalization: Cloud-device Collaborative Data Augmentation for Efficient On-device Language Model

### v1

@@ 1. Introduction
Recently, Large Language Models (LLMs) have become a cornerstone of contemporary Natural Language Processing (NLP) research and industry applications due to their exceptional abilities in text understanding and generation (Radford and Narasimhan, 2018; Ray, 2023; Naveed et al., 2024). These models have achieved remarkable success and transformed numerous areas of NLP, such as translation, summarization, and conversational AI (Thirunavukarasu et al., 2023; Hu et al., 2024a; Wang et al., 2024).
@@
Despite their advancements, existing LLMs face two significant limitations that hinder their broader adoption: (1) Lack of Personalization. LLMs are designed as universal models, which limits their ability to generate responses tailored to users’ personalized preferences and interests; (2) Dependence on Cloud Infrastructure. The powerful LLMs are typically trained and deployed on cloud servers due to their high computational demands. This architecture not only relies on stable and high-speed network connections to transmit user queries and deliver responses, but also requires a long time for LLM inference. However, these conditions are often unmet in real-world scenarios, particularly on mobile platforms such as smartphones and smart vehicles. For instance, LLM-based assistants are increasingly integrated into smart vehicle systems, but these vehicles frequently experience inconsistent network connectivity due to their mobility. In remote regions with weak or no network coverage, such cloud-based LLM services could become entirely inaccessible as the system goes offline. In addition, considering the time-sensitive applications, an LLM-powered service will lead to a significant inference latency overhead in reality, limiting the feasibility of the service. Consequently, there is a growing need for personalized LLMs that can run directly on user devices. Such models must address user-specific needs while operating efficiently on edge devices, free from the constraints of cloud connectivity.
@@
Some recent efforts have explored techniques for enabling personalization in LLMs, which can be generally categorized into prompt-based methods and fine-tuning-based methods. The prompt-based approaches format personalized prompts to leverage the in-context learning capabilities of LLMs. That is to say, all users share the same model, but personalized prompts are used to guide the generation process. For example, Christakopoulou et al. (2023) incorporates users’ historical data into prompts to enhance generation performance. And to conquer the input length limitation when users’ historical data are too long, some research employs retrieval-augmentation generation (RAG) to augment user’s query by adding the most relevant history information into prompt (Richardson et al., 2023; Salemi et al., 2023; Li et al., 2024b, c). On the other hand, fine-tuning methods directly optimize the parameters of LLMs to adapt to users’ personalized data distributions (Tan et al., 2024b, a; Park et al., 2024; Li et al., 2024c; Zhuang et al., 2024). However, these approaches face significant scalability issues, as the cloud server must fine-tune a separate model for each user. As the number of users grows, centralized computation becomes a major bottleneck, making the fine-tuning-based personalization method impractical.
@@
In this paper, we take a first step toward developing a learning framework for personalized on-device language models (LMs). Our approach leverages cloud-based large language models (LLMs) to address the challenge of personal data scarcity and introduces a cloud-device collaborative framework to ensure scalability. Specifically, the cloud-based LLM generates synthetic data tailored to each user’s limited local data, thereby augmenting the user’s dataset and transferring relevant knowledge from the cloud LLM to the on-device LM. Once users receive the synthetic data, we apply parameter-efficient fine-tuning (PEFT) techniques to optimize their on-device LMs using both the synthetic and local personal data entirely on the user’s device. This decentralized training strategy avoids the scalability bottleneck of requiring the cloud server to fine-tune models for all users. After training, the personalized on-device LM is deployed locally, allowing inference to be performed without network connectivity, which reduces both network dependency and inference latency. To evaluate the framework’s effectiveness, we conduct extensive experiments on public datasets, and experimental results demonstrate that the proposed method can achieve promising performance in personalized classification and generation tasks.
@@ 2.1. Personalization of LLMs
Fine-tuning Based Personalization. Parameter-efficient fine-tuning (PEFT) offers an effective way to optimize LLMs for users’ personal distributions by modifying only a small subset of parameters (Hu et al., 2021; Dettmers et al., 2023). For example, OPPU proposed a PEFT-based personalized LLM, which fine-tunes the LoRA adapter on user profiles for each user, to store user knowledge on PEFT parameters (Tan et al., 2024b). Building on this work, PER-PCS aggregates fine-tuned LoRA adapters from multiple users into a shared adapter pool, which can be leveraged to generate a personalized LLM for a target user by merging multiple LoRA adapters (Tan et al., 2024a). Reinforcement learning is also applied with PEFT to achieve better performance(Cheng et al., 2023; Li et al., 2024c; Park et al., 2024).
@@ 2.3. Textual Data Augmentation
Recent progress in LLMs has shown superiority in producing high-quality, contextually relevant augmented text by analysing and reconstructing sentences to enhance both diversity and richness. For example, AugGPT (Dai et al., 2025) leverages the ChatGPT model to rewrite sentences while preserving dataset consistency, and Self-LLMDA (Li et al., 2024a) automatically generates and selects optimal instructions to guide LLMs in producing augmented samples. Meanwhile, several studies (Lu and Lam, 2023; Li et al., 2023; Chen et al., 2024; Hu et al., 2024b; Zhao et al., 2024) have explored paraphrasing-based augmentation with LLMs, where the semantic content is preserved while surface forms, such as wording and structure, are altered. Such methods are particularly beneficial for scenarios with scarce data, as they need to enrich data without altering semantic meaning. However, these prompting-based techniques are not directly aligned with our objective of generating personalized text that reflects individual users’ preferences.
@@ 3. Research Problem Formulation
$t_{u}$ The history happened before query time t of user $u$.
@@
This paper explores fine-tuning personalized on-device language models (LMs), incorporating two key concepts: personalized LMs and on-device LMs. Unlike general LMs, which produce output sequences solely based on the input sequence, a personalized LM generates responses by considering both the user’s query $x$ and their profile $D_{u}$. We define the user profile as a collection of the user’s historical input-output pairs: i.e., $D_{u}=\{(x_{u1},y_{u1}),(x_{u2},y_{u2}),\dots,(x_{ut_{u}},y_{ut_{u}})\}$, where $t_{u}$ indicates the history before query time $t$.
@@ 4. Proposed Method
[…]
Overview of framework includes five steps. user data uploading, data augmentation with server LLM, synthetic data selection, synthetic data downloading, and on-device LM personalization fine-tuning.
@@ 4.2. Data Augmentation with Server LLM
[…]
Using prompts to generate inputs, and passing synthetic inputs to generate outputs.
@@ 4.3. Synthetic Data Selection
Filter 1: Semantic consistency filter. Reliable synthetic data should preserve the semantics of the original statement without introducing hallucinated content. Natural Language Inference (NLI) models are trained to determine whether a given “hypothesis” and “premise” entail, contradict, or are neutral to each other. Therefore, we employ a small NLI model $M_{NLI}$ (Liu et al., 2022) as the semantic evaluator, which provides a semantic consistent score between the synthetic and original samples:
@@ 5.1. Experimental Settings
To promote the personalization phenomenon, we designate two groups of target users: the 100 most active users with the longest history logs and an additional 100 users randomly sampled from the dataset, while using all remaining users for base model training. Our objective is to obtain an on-device personalized LM for each user among these 200 users. Sampling strategies are widely adopted to reduce the computational overhead of personalization evaluation, as demonstrated in many existing works evaluated on the LaMP benchmark. For instance, HYDRA (Zhuang et al., 2024) randomly selected 100 users for training and an additional 50 users for evaluation, while FERMI (Kim and Yang, 2025) selected 50 users for training and evaluation. Similarly, OPPU (Tan et al., 2024b), Personaized Piece (Tan et al., 2024a), and PriME (Kim et al., 2025) selected the 100 most active users with the longest history logs for training and evaluation. In this paper, we not only follow the same sampling strategies of prior works (Tan et al., 2024b, a; Kim et al., 2025), but also adopt a random sampling approach, which offers a statistically representative assessment of overall dataset performance.
@@
Following LaMP (Salemi et al., 2023), we use accuracy and F1-score for the LaMP-1 and LaMP-2, mean absolute error (MAE) and root mean square error (RMSE) for the LaMP-3, and ROUGE-1, ROUGE-L, and BERTScore-F1 (BERT-F1) for LaMP-4, LaMP-5, and LaMP-7. Except for MAE and RMSE, where lower values are better, all other metrics with higher values indicate better performance. Accuracy measures the proportion of correct predictions among the total number of cases processed:
(10) $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN},$
@@
By default, we set the LLM generation samples $k$ to 5 in all experiments. We apply the LoRA adapter on all linear layers of the on-device model, and set the LoRA rank $r$ to 16 and scaling factor $\alpha$ to 8. We quantize the on-device model weight in the NF4 data type and use bfloat16 for computation. Followed by the Qwen2.5 technique report (Yang et al., 2024), we used the multinomial sampling decoding with temperature $\tau_{temp}=0.7$ to balance the computational efficiency and sampling diversity of data generation. We implement all the experiments using Pytorch (Paszke et al., 2019) and HuggingFace library (Wolf et al., 2020) on an NVIDIA RTX A5000 GPU.
@@ 5.3. Ablation Study (RQ2)
[…]
The performance increases as the number of LLM-generated samples increasing from 0 to 5.
@@ 5.5. Efficiency Analysis (RQ4)
Models
Storage
Size (GB)
Workstation Android
@@ 5.6. Generalization of Framework (RQ5)
[…]
On all 6 tasks, a user’s real profile is more similar to their own synthetic profile.
@@ 5.8. Case Study
[…]
A case study for user 11003279. RAG personalized achieves 0.09 ROUGE-1 score. Direct fine-tune achieves 0.07. And our method achieves 0.43.
@@ 6. Limitations and Ethical Considerations
Several limitations are concerned with our work. Firstly, due to dataset constraints, our study aims to deploy a personalized model to generate responses on a specific task for each user, ignoring the user behaviors from other tasks and domains. For example, for the user who engages in news headline generation and scholarly title generation tasks, both tasks could provide the user’s stylistic pattern preference. Nevertheless, in the future, we believe CDCDA-PLM can be applied to any NLP task across different domains. Secondly, the data quality of LLM augmentation can be affected by the cloud-based LLM. Exploring a larger LLM or multiple LLMs to augment user data remains an area for future investigation.
Training a personalized model heavily relies on personal data, which may leak sensitive or private information of users. Sharing user data with server LLM for user personal data augmentation also leads to privacy concerns. Therefore, it is important to investigate further robust methods for privacy protection in cloud-server LLM data augmentation. In addition, a personalized model aims to generate content aligning with user preferences and interests shown in user data. However, personalization models may be trained with user data consisting of biased and unfair information, leading to harmful responses. Within CDCDA-PLM, the biased data is uploaded to server LLM for augmentation, which further negatively affects the on-device model. Future works may explore strategies to avoid sharing or augmenting harmful data on the server LLM.
@@ Acknowledgements.
The Australian Research Council partially supports this work under the streams of Future Fellowship (Grant No. FT210100624), the Discovery Project (Grant No. DP240101108), and the Linkage Projects (Grant No. LP230200892 and LP240200546).

### v2

@@ 1. Introduction
Recently, Large Language Models (LLMs) have become a cornerstone of contemporary Natural Language Processing (NLP) research and industry applications due to their exceptional abilities in text understanding and generation (Radford and Narasimhan, 2018; Ray, 2023; Naveed et al., 2024). These models have achieved remarkable success and transformed numerous areas of NLP, such as translation, summarization, and conversational AI (Thirunavukarasu et al., 2023; Hu et al., 2024b; Wang et al., 2024).
@@
Despite their advancements, existing LLMs face two significant limitations that hinder their broader adoption: (1) Lack of Personalization. LLMs are designed as universal models, which limits their ability to generate responses tailored to users’ personalized preferences and interests; (2) Dependence on Cloud Infrastructure. The powerful LLMs are typically trained and deployed on cloud servers due to their high computational demands. This architecture not only relies on stable and high-speed network connections to transmit user queries and deliver responses, but also requires a long time for LLM inference. However, these conditions are often unmet in real-world scenarios, particularly on mobile platforms such as smartphones and smart vehicles (Xia et al., 2022; Yin et al., 2025). For instance, LLM-based assistants are increasingly integrated into smart vehicle systems, but these vehicles frequently experience inconsistent network connectivity due to their mobility. In remote regions with weak or no network coverage, such cloud-based LLM services could become entirely inaccessible as the system goes offline. In addition, considering the time-sensitive applications, an LLM-powered service will lead to a significant inference latency overhead in reality, limiting the feasibility of the service. Consequently, there is a growing need for personalized LLMs that can run directly on user devices. Such models must address user-specific needs while operating efficiently on edge devices, free from the constraints of cloud connectivity (Wang et al., 2020; Zheng et al., 2024b).
@@
Some recent efforts have explored techniques for enabling personalization in LLMs, which can be generally categorized into prompt-based methods and fine-tuning-based methods. The prompt-based approaches format personalized prompts to leverage the in-context learning capabilities of LLMs. That is to say, all users share the same model, but personalized prompts are used to guide the generation process. For example, Christakopoulou et al. (2023) incorporates users’ historical data into prompts to enhance generation performance. And to conquer the input length limitation when users’ historical data are too long, some research employs retrieval-augmentation generation (RAG) to augment user’s query by adding the most relevant history information into prompt (Richardson et al., 2023; Salemi et al., 2023; Li et al., 2024a, b). On the other hand, fine-tuning methods directly optimize the parameters of LLMs to adapt to users’ personalized data distributions (Tan et al., 2024b, a; Park et al., 2024; Li et al., 2024b; Zhuang et al., 2024). However, as the number of users grows, these approaches face significant scalability challenges because the server’s computational cost increases at least linearly with the user base, making large-scale deployment impractical (Long et al., 2024; Zheng et al., 2024a). Specifically, since a separate personalized model must be trained for each user, a server-based design would require training N models for N users. As the number of users grows, this results in a significant computational burden on the server, making the fine-tuning-based personalization method impractical (Xia et al., 2023a, b; Yuan et al., 2025).
@@
In this paper, we take a first step toward developing a learning framework for personalized on-device language models (LMs). Our approach leverages cloud-based large language models (LLMs) to address the challenge of personal data scarcity and introduces a cloud-device collaborative framework to ensure scalability. Specifically, the cloud-based LLM generates synthetic data tailored to each user’s limited local data, thereby augmenting the user’s dataset and transferring relevant knowledge from the cloud LLM to the on-device LM. Once users receive the synthetic data, we apply parameter-efficient fine-tuning (PEFT) techniques to optimize their on-device LMs using both the synthetic and local personal data entirely on the user’s device. This decentralized training strategy avoids the scalability bottleneck of requiring the cloud server to fine-tune models for all users, parallelized personalization across a large user base. After training, the personalized on-device LM is deployed locally, allowing inference to be performed without network connectivity, which reduces both network dependency and inference latency. To evaluate the framework’s effectiveness, we conduct extensive experiments on public datasets, and experimental results demonstrate that the proposed method can achieve promising performance in personalized classification and generation tasks.
@@ 2.1. Personalization of LLMs
Fine-tuning Based Personalization. Parameter-efficient fine-tuning (PEFT) offers an effective way to optimize LLMs for users’ personal distributions by modifying only a small subset of parameters (Hu et al., 2021; Dettmers et al., 2023). For example, OPPU proposed a PEFT-based personalized LLM, which fine-tunes the LoRA adapter on user profiles for each user, to store user knowledge on PEFT parameters (Tan et al., 2024b). Building on this work, PER-PCS aggregates fine-tuned LoRA adapters from multiple users into a shared adapter pool, which can be leveraged to generate a personalized LLM for a target user by merging multiple LoRA adapters (Tan et al., 2024a). Reinforcement learning is also applied with PEFT to achieve better performance(Cheng et al., 2023; Li et al., 2024b; Park et al., 2024).
@@ 2.3. Textual Data Augmentation
Recent progress in LLMs has shown superiority in producing high-quality, contextually relevant augmented text by analysing and reconstructing sentences to enhance both diversity and richness. For example, AugGPT (Dai et al., 2025) leverages the ChatGPT model to rewrite sentences while preserving dataset consistency, and Self-LLMDA (Li et al., 2024c) automatically generates and selects optimal instructions to guide LLMs in producing augmented samples. Meanwhile, several studies (Lu and Lam, 2023; Li et al., 2023; Chen et al., 2024; Hu et al., 2024a; Zhao et al., 2024) have explored paraphrasing-based augmentation with LLMs, where the semantic content is preserved while surface forms, such as wording and structure, are altered. Such methods are particularly beneficial for scenarios with scarce data, as they need to enrich data without altering semantic meaning. However, these prompting-based techniques are not directly aligned with our objective of generating personalized text that reflects individual users’ preferences.
@@ 3. Research Problem Formulation
$t$ The query time of a user.
@@
This paper explores fine-tuning personalized on-device language models (LMs), incorporating two key concepts: personalized LMs and on-device LMs. Unlike general LMs, which produce output sequences solely based on the input sequence, a personalized LM generates responses $y_{u,t}$ by considering both the user’s query $x_{u,t}$ and their profile $D_{u}$. We define the user profile as a collection of the user’s historical input-output pairs: i.e., $D_{u}=\{(x_{u,1},y_{u,1}),(x_{u,2},y_{u,2}),\dots,(x_{u,t-1},y_{u,t-1})\}$, which represent the user’s interactions that occurred prior to the query time $t$.
@@ 4. Proposed Method
Refer to caption Overview of framework includes five steps. user data uploading, data augmentation with server LLM, synthetic data selection, synthetic data downloading, and on-device LM personalization fine-tuning.
@@ 4.2. Data Augmentation with Server LLM
Refer to caption Using prompts to generate inputs, and passing synthetic inputs to generate outputs.
@@ 4.3. Synthetic Data Selection
Filter 1: Semantic consistency filter. Reliable synthetic data should preserve the semantics of the original statement without introducing hallucinated content. Natural Language Inference (NLI) models are trained to determine whether a “hypothesis” preserves the similar meaning as the “premise”. Specifically, the NLI model classified the semantic relationship into one of three categories: (1) Entailment: The meaning of the hypothesis can be logically inferred from the premise. (2) Contradiction: The meaning of the hypothesis directly contradicts the premise. (3) Neutral: There is no clear logical connection or conflict between the premise and the hypothesis. For each given pairs $(x,x_{syn})$, semantic evaluator outputs an entailment probability in both direction, $M_{NLI}(x\Rightarrow x_{syn})$ and $M_{NLI}(x_{syn}\Rightarrow x)$. We treat the sample as semantically consistent if both probabilities exceed a threshold $\epsilon_{scf}$. Therefore, we employ a small NLI model $M_{NLI}$ (Liu et al., 2022) as the semantic evaluator and define the filters as:
@@ 5.1. Experimental Settings
To promote the personalization phenomenon, we designate two groups of target users: the 100 most active users with the longest history logs and an additional 100 users randomly sampled from the dataset, while using all remaining users for base model training. Our objective is to obtain an on-device personalized LM for each user among these 200 users. Sampling strategies are widely adopted to reduce the computational overhead of personalization evaluation, as demonstrated in many existing works evaluated on the LaMP benchmark. For instance, HYDRA (Zhuang et al., 2024) randomly selected 100 users for training and an additional 50 users for evaluation, while FERMI (Kim and Yang, 2025) selected 50 users for training and evaluation. Similarly, OPPU (Tan et al., 2024b), Personaized Piece (Tan et al., 2024a), and PriME (Kim et al., 2025) selected the 100 most active users with the longest history logs for training and evaluation. In real-world deployment, these on-device personalized LMs can be executed in parallel and independently on each user’s device. However, for our evaluation experiments, we have to build these user-specific personalized models one by one due to the equipment limitation, leading to a relatively long evaluation time. Therefore, we have to sample only a part of the users for evaluation. We believe this may also be the main reason that previous works use sampling user strategies rather than full-set users for evaluation. In this paper, we not only follow the same sampling strategies of prior works (Tan et al., 2024b, a; Kim et al., 2025), but also adopt a random sampling approach, which offers a statistically representative assessment of overall dataset performance.
@@
Following LaMP (Salemi et al., 2023), we use accuracy and F1-score for the LaMP-1 and LaMP-2, mean absolute error (MAE) and root mean square error (RMSE) for the LaMP-3, and ROUGE-1, ROUGE-L, and BERTScore-F1 (BERT-F1) for LaMP-4, LaMP-5, and LaMP-7. Except for MAE and RMSE, where lower values are better, all other metrics with higher values indicate better performance. The mathematical equations of metrics are presented in Appendix B
@@
By default, we set the LLM generation samples $k$ to 5 in all experiments. For each case, we independently execute the model five times with different random seeds and report the averaged outcomes, and all results are statistically significant at $p<$ 0.05. We apply the LoRA adapter on all linear layers of the on-device model, and set the LoRA rank $r$ to 16 and scaling factor $\alpha$ to 8. We quantize the on-device model weight in the NF4 data type and use bfloat16 for computation. Followed by the Qwen2.5 technique report (Yang et al., 2024), we used the multinomial sampling decoding with temperature $\tau_{temp}=0.7$ to balance the computational efficiency and sampling diversity of data generation. We implement all the experiments using Pytorch (Paszke et al., 2019) and HuggingFace library (Wolf et al., 2020) on an NVIDIA RTX A5000 GPU.
@@ 5.3. Ablation Study (RQ2)
Refer to caption The performance increases as the number of LLM-generated samples increasing from 0 to 5.
@@ 5.5. Efficiency Analysis (RQ4)
Models Storage Size (GB) Workstation Android
@@ 5.6. Generalization of Framework (RQ5)
Refer to caption On all 6 tasks, a user’s real profile is more similar to their own synthetic profile.
@@ 5.8. Case Study
Refer to caption A case study for user 11003279. RAG personalized achieves 0.09 ROUGE-1 score. Direct fine-tune achieves 0.07. And our method achieves 0.43.
@@ 6. Limitations and Ethical Considerations
Several limitations are concerned with our work. Firstly, due to dataset constraints, our study aims to deploy a personalized model to generate responses on a specific task for each user, ignoring the user behaviors from other tasks and domains. For example, for the user who engages in news headline generation and scholarly title generation tasks, both tasks could provide the user’s stylistic pattern preference. Moreover, in the future, we believe CDCDA-PLM can be applied to any NLP task across different domains. Secondly, the data quality of LLM augmentation can be affected by the cloud-based LLM. Exploring a larger LLM or multiple LLMs to augment user data remains an area for future investigation. Thirdly, although CDCDA-PLM relies on the cloud LLM only during the synthetic data generation stage, it may limit applicability to the users in offline-first or highly restricted settings. To address this challenge, our framework can be extended to employ a lightweight local LLM as on-device synthetic data generator.
CDCDA-PLM uploading user data to server LLM for cloud-based data augmentation may leads to privacy concerns. Therefore, our framework assumes that user data can be utilized under appropriate legal and ethical agreements, as is common in many real-world personalization systems (e.g., recommender services or cloud-based assistants). These agreements typically allow data collection and processing for the explicit purpose of improving user experience, often with user consent and privacy safeguards in place. Nevertheless, CDCDA-PLM can be extended with additional privacy-preserving strategies to further reduce potential risks. First, data minimization can be applied by transmitting only compact user profile summaries instead of raw text, thereby preventing the server from directly accessing sensitive content. Secondly, differential privacy (DP) can be incorporated during user profile uploading by adding controlled noise, which prevents the server from reconstructing user history. Therefore, it is important to investigate further robust methods for privacy protection in cloud-server LLM data augmentation.
In addition, a personalized model aims to generate content aligning with user preferences and interests shown in user data. However, personalization models may be trained with user data consisting of biased and unfair information, leading to harmful responses. Within CDCDA-PLM, the biased data is uploaded to server LLM for augmentation, which further negatively affects the on-device model. Future works may explore strategies to avoid sharing or augmenting harmful data on the server LLM.
@@ Acknowledgements.
The Australian Research Council partially supports this work under the streams of Future Fellowship (Grant No. FT210100624), the Discovery Project (Grant No. DP240101108 and DP260100326), and the Linkage Projects (Grant No. LP230200892 and LP240200546).
@@ A.6. LaMP-7: Personalized Tweet Paraphrasing
Appendix B Equations of Evaluation Metrics
Accuracy measures the proportion of correct predictions among the total number of cases processed:
(10) $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN},$